OpenStudy (anonymous):

What is the solution of the equation? b+6/5 = 10 A. 44 B. –4 C. 56 D. 8

OpenStudy (anonymous):

\[\frac{b+6}{5}=10\] multiply both sides by 5 to get \[b+6=50\] last step is to subtract \(6\) from both sides

OpenStudy (anonymous):

i dont know how to make the signs for the problem

OpenStudy (anonymous):

im still not understanding can you explain further please

OpenStudy (anonymous):

-4??